OpenStudy (anonymous):

A car burns 420. g of N2, according to the equation N2 + 2O2 -> 2NO2 3NO2 + H2O -> 2HNO3 + NO How many grams of HNO3 will be produced?

OpenStudy (frostbite):

Calculate the amount of substance for nitrogen, assume that the reactions are irreversible and make sure the stoichiometry is right and calculate the amount of substance for nitric acid then just use the relation between amount of substance and molecular mass to calculate the mass.

OpenStudy (anonymous):

So, I got

OpenStudy (frostbite):

\[n(N _{2})=\frac{ m(N _{2} }{ M(N _{2}) }\] \[n(NO _{2})=2*n(N _{2})\] \[\frac{ n(HNO _{3}) }{ 2 }=\frac{ n(NO _{2}) }{ 3 } \rightarrow n(HNO _{3})=\frac{ 2 }{ 3 } n(NO _{2})\] \[m(NHO _{3})=n(NHO _{3})M(NHO _{3})\] What I can see.

OpenStudy (anonymous):

Oh, you have to go all the way through the equation? I didn't think of having to use NO2.

OpenStudy (frostbite):

well it is a 100% sure way.

OpenStudy (frostbite):

So longer you remember the stoichiometry your way might work too... haven't checked.

OpenStudy (anonymous):

No, you were right! I had to go all through the equation. Thank you! :)

OpenStudy (frostbite):

But if you substitute everything together into 1 single equation is up to you, or you can calculate the things part by part. :)

OpenStudy (frostbite):

And no problem.

OpenStudy (anonymous):

You are a genius. Want to help me with a couple more? :)

OpenStudy (frostbite):

No I'm just me :) and unfortunately I can't... I should be in bed 15 min ago. but tag one called "aaronq" he know all this as well.

OpenStudy (anonymous):

Well thank you again! It's greatly appreciated. :)

OpenStudy (frostbite):

No problem at all.